OpenStudy (anonymous):

how to factor the following trinomials forms: x^2-7x+12

Parth (parthkohli):

-3 - 4 = -7 -3 * -4 = 12\[\rm x^2 - 3x - 4x + 12\]By grouping first two and last two, you get,\[\rm x(x - 3) - 4(x - 3)\]

OpenStudy (anonymous):

Sum = -7 Product = 12 Find two Such numbers

OpenStudy (anonymous):

so how would i explain this?

OpenStudy (anonymous):

-3 and -4 nw split -7x = -3x - 4x x^2-7x+12 x^2-3x-4x+12 x(x-3) -4(x-3) (x-3) (x-4)